# From Fuzzy to Exact: The Halo Architecture for Infinite-Depth Reasoning via Rational Arithmetic

### v3

@@ Abstract
The pursuit of scale in deep learning has entrenched a trade-off: computational throughput is prioritized at the expense of numerical precision. We argue this compromise is fundamentally at odds with the requirements of general intelligence. We propose the Exactness Hypothesis: high-order causal reasoning—a cornerstone of AGI—demands a substrate supporting arbitrary-precision, logically consistent arithmetic. We trace prevalent LLM failures, such as logical hallucinations and incoherence, to the inherent limitations of IEEE 754 floating-point arithmetic, where approximation errors compound catastrophically in deep functions.
As a solution, we present the Halo Architecture, which transitions the computational foundation from approximate reals ($\mathbb{R}$) to exact rationals ($\mathbb{Q}$). Halo is realized through a custom Exact Inference Unit (EIU), whose design—featuring asynchronous MIMD reduction and dual-modular redundancy—resolves the performance and reliability bottlenecks of exact computation at scale. Crucially, we theoretically prove that Halo’s number-theoretic quantization yields a quadratic error decay ($\mathcal{O}(D^{-2})$), strictly superior to the linear barrier of standard fixed-point arithmetic.
In rigorous simulations, 600B-parameter BF16 models fail in chaotic systems within steps, while Halo sustains perfect numerical fidelity indefinitely. Our work posits exact arithmetic as non-negotiable for advancing reasoning-capable AGI and provides a co-designed hardware-software path toward verifiable, exascale-ready AI systems.
@@ 1 Introduction
We argue that this is due to a fundamental mismatch between the physical substrate (approximate floating-point hardware) and the mathematical requirements of rigorous logic. Logic is discrete and precise; floating-point arithmetic is continuous and approximate. We observe that in deep recursive inference, numerical noise $\epsilon$ accumulates exponentially ($\mathcal{O}(e^{L})$), eventually drowning out the causal signal. We term this phenomenon Semantic Drift. This aligns with recent findings by Thinking Machines Lab [Thinking Machines Lab, 2025], who pinpointed floating-point non-associativity as the culprit for inference instability.
@@
However, Halo goes beyond mere exactness. We introduce the Ring Mechanism to solve the tractability of rational arithmetic. Unlike standard quantization which truncates values linearly ($\mathcal{O}(D^{-1})$), our approach utilizes Diophantine Approximation to project states onto the Farey Sequence. We theoretically prove that this achieves a quadratic error decay ($\mathcal{O}(D^{-2})$), allowing Halo to preserve significantly higher information density per bit. We term this process Semantic Crystallization: using the discrete nature of number theory to filter out numerical noise while locking in the logical signal.
@@ 4.1 Algorithmic Formulation
To address the precision-efficiency trade-off, Halo partitions the computational flow into two distinct regimes. We formally define these core components:
  • The Light (Exact Stream): The primary datapath where calculations occur in the arbitrary-precision rational field $\mathbb{Q}$ (represented as exact fractions $p/q$). In this stream, operations are chemically pure: no truncation or rounding occurs, ensuring $\epsilon=0$. Unlike floating-point numbers which drift, "The Light" ensures inference depth is limited only by time, not precision.
  • The Ring (Periodic Reset): A discrete bottleneck mechanism. It acts as a "Gatekeeper of Complexity," projecting the potentially complex rational state onto a fixed, low-bitwidth lattice to prevent the "explosion" of denominator complexity.
4.2 Halo-V: Vector-Symbolic Alignment (Baseline)
In our initial formulation (Halo-V), we implemented The Ring as a Vector-Symbolic Quantizer. This approach aligns the rational state with a statistically learned "Language of Thought" using VQ-VAE [Van Den Oord et al., 2017]. Algorithms 1 and 2 detail this baseline process, where the reset is triggered by a fixed schedule ($K$) and relies on learned Encoders/Decoders.
Algorithm 1 Halo-V: End-to-End Inference (Fixed Schedule)
0: Input token sequence $X\in\mathbb{V}^{L}$
0: Recurrent Depth $T$ (Thinking Steps)
0: Ring Interval $K$ (Grounding Frequency)
0: Next token probability distribution $P(y|X)$
1:Stage 1: Prelude
@@ 4.2 Halo-V: Vector-Symbolic Alignment (Baseline)
4:Stage 2: The Light (Recurrent Core)
5:for $t=1$ to $T$ do
6:  $H_{attn}\leftarrow\text{RationalAttention}(H_{t-1})$
7:  $H_{mlp}\leftarrow\text{RationalFFN}(H_{attn})$
8:  $H_{temp}\leftarrow H_{mlp}+H_{t-1}$
9:  Stage 3: The Ring (Fixed Schedule)
10:  if $t\mod K==0$ then
11:   {Vector-Symbolic Collapse}
12:   $H_{t}\leftarrow\text{TheRing}(H_{temp})$
13:  else
14:   $H_{t}\leftarrow H_{temp}$
15:  end if
16:end for
17:Stage 4: Coda
@@
[…]
20:return $\text{Softmax}(logits)$
Algorithm 2 Halo-V Ring Mechanism (Learned Quantizer)
1:Function TheRing($H_{\mathbb{Q}}$)
2: {1. Projection to Latent Manifold (Floating Point)}
[…]
4: {2. Semantic Auto-Encoding (Denoising)}
[…]
7: {3. Re-grounding to Rational Space}
[…]
10:return $H_{reset}$
11:End Function
12:Function RationalAttention($Q,K,V$)
[…]
15:return $A\cdot V$
16:End Function
@@ 4.3 Implementation: The Exact Inference Unit (EIU)
    Interrupt-Driven Registers: EIU registers are designed to operate in a free-running mode with dynamic bit-width. Crucially, they possess a Hardware Saturation Flag that triggers an interrupt strictly when the denominator’s MSB reaches the capacity limit (e.g., 4096 bits).
@@
  4. 4.
    Dual-Modular Redundancy: To prevent Silent Data Corruption (SDC), operations are verified modulo Mersenne primes (see Appendix C).
4.4 Halo-N: Number-Theoretic Quantization via Diophantine Approximation
While Halo-V provides a robust logical baseline, the capabilities of the EIU enable a more advanced paradigm: Halo-N (Number-Theoretic). By leveraging the specific hardware features described above, Halo-N evolves the architecture from "Heuristic Approximation" to "Mathematical Optimality."
1. From Fixed Schedule to Natural Saturation
In Halo-V, the Ring is triggered at a fixed interval $K$. This is rigid. Halo-N utilizes the EIU’s Interrupt-Driven Registers to implement a "Natural Saturation" policy. The Ring reset is triggered only when the hardware flag asserts overflow. This transforms $K$ from a hyperparameter into a dynamic, physics-driven variable.
2. From VQ-VAE to Diophantine Solver
In Halo-V, The Ring relies on a learned Encoder/Decoder. Halo-N utilizes the EIU’s Asynchronous GCD Engines to replace this with a Diophantine Solver. Instead of learning a codebook, we project the state onto the Farey Sequence $\mathcal{F}_{N}$ using Coordinate-wise Diophantine Approximation. The mathematical process of finding the best rational approximation via Continued Fractions is algorithmically isomorphic to the Euclidean GCD algorithm, allowing the EIU to perform "Semantic Crystallization" without requiring any learned parameters.
Algorithms 3 and 4 detail this evolved process.
Algorithm 3 Halo-N: End-to-End Inference (Adaptive Saturation)
0: Input token sequence $X\in\mathbb{V}^{L}$
0: Recurrent Depth $T$ (Thinking Steps)
0: Saturation Limit $B_{max}$ (e.g., 4096 bits)
0: Next token probability distribution $P(y|X)$
1:Stage 1: Prelude
[…]
4:Stage 2: The Light (Recurrent Core)
5:for $t=1$ to $T$ do
6:  $H_{attn}\leftarrow\text{RationalAttention}(H_{t-1})$
7:  $H_{mlp}\leftarrow\text{RationalFFN}(H_{attn})$
8:  $H_{temp}\leftarrow H_{mlp}+H_{t-1}$
9:  Stage 3: The Ring (Interrupt-Driven Reset)
10:  if $\text{CheckSaturation}(H_{temp},B_{max})$ then
11:   {Triggered by Hardware Overflow Flag}
12:   {Diophantine Projection}
13:   $H_{t}\leftarrow\text{TheRing}(H_{temp})$
14:  else
15:   $H_{t}\leftarrow H_{temp}$
16:  end if
17:end for
18:Stage 4: Coda
[…]
21:return $\text{Softmax}(logits)$
Algorithm 4 Halo-N Ring Mechanism (Diophantine Solver)
1:Function TheRing($H_{\mathbb{Q}}$)
2: {Projects state onto the Cartesian Farey Lattice $\mathcal{F}^{d}$}
[…]
4: {1. Coordinate-wise Decomposition (MIMD)}
5:for all element $h_{ij}$ in tensor $H_{\mathbb{Q}}$ do
6:  $x_{val}\leftarrow h_{ij}.num/h_{ij}.denom$
7:  {2. Best Rational Approximation (Continued Fraction)}
8:  {Find $p/q$ minimizing $|x_{val}-p/q|$ s.t. $q\leq D_{max}$}
9:  $p,q\leftarrow\text{TruncatedEuclidean}(h_{ij}.num,h_{ij}.denom,D_{max})$
10:  $H_{new}[i,j]\leftarrow\text{Rational}(p,q)$
11:end for
12:return $H_{new}$
13:End Function
14:Function RationalAttention($Q,K,V$)
[…]
17:return $A\cdot V$
18:End Function
4.5 Training Objective: The Rational Alignment Loss
To ensure that the continuous representations learned by the network are amenable to lossless rational projection, we introduce a novel auxiliary loss term: the Rational Alignment Loss ($\mathcal{L}_{ring}$).
Unlike standard Vector Quantization (VQ) which maps vectors to an arbitrary learned codebook [Van Den Oord et al., 2017], Halo maps vectors to the fixed algebraic field of logical rationals $\mathbb{Q}_{K}=\{\frac{n}{d}\mid|d|\leq K\}$.
Let $z_{e}(x)$ be the continuous embedding (Float) and $z_{q}(x)$ be the projected rational approximation (The Ring output). We define the objective as:
[…]
where $\mathcal{L}_{NTP}$ is the Next Token Prediction loss, and $\text{sg}[\cdot]$ denotes the stop-gradient operator.
  • The first term is the Commitment Loss, forcing the continuous encoder to commit to a specific rational value.
  • The second term optimizes the rational projection (if $z_{q}$ is differentiable via relaxation) or acts as a proxy for codebook alignment.
Significance for Halo-N: While derived from the VQ-VAE paradigm used in Halo-V, this objective is equally critical for Halo-N to minimize the “Diophantine Error” during hardware resets. By strictly constraining the target space to $\mathbb{Q}_{K}$ (where denominators are bounded by $K$), $\mathcal{L}_{ring}$ acts as a complexity penalty. It actively discourages the network from generating high-entropy states that would require massive denominators to represent, effectively “pulling” potentially long, irreducible rational outputs towards lower-complexity, shorter-length rationals (e.g., simple fractions) that fit within the hardware’s optimal precision window.
4.6 Transcending the Irrational: Rational Approximations
A core challenge in deploying Rational Arithmetic for Deep Learning is that standard activation functions (e.g., GeLU, Softmax) and normalization layers (LayerNorm) rely on transcendental operations ($e^{x}$, $\sqrt{x}$) that yield irrational numbers. To maintain the "Logical Rigidity" of the Halo architecture, we cannot simply cast these to float approximations. Instead, we implement them as Convergent Rational Series within the Exact Inference Unit (EIU).
Rational Softmax via Constructive Definitions. The standard attention mechanism requires computing $A=\text{softmax}(QK^{T}/\sqrt{d})$. We first replace the scaling factor $1/\sqrt{d}$ with an integer shift $2^{-k}$. For the exponential function, we strictly adhere to its mathematical definition as the limit of an infinite rational series. For any $x\in\mathbb{Q}$, the partial sum constitutes a dense rational approximation:
$\text{RatExp}(x,N)=\sum_{k=0}^{N}\frac{x^{k}}{k!}\xrightarrow{N\to\infty}e^{x}$ (9)
where $N$ is a dynamic precision parameter determined by the EIU. Crucially, this expansion is not merely an approximation technique but the constructive definition of the exponential function itself. Unlike floating-point implementations which rely on hardware-specific lookup tables with fixed error, RatExp allows us to trade compute time for arbitrary precision. The Rational Softmax is then defined as:
$\text{RatSoftmax}(z)_{i}=\frac{\text{RatExp}(z_{i},N)}{\sum_{j}\text{RatExp}(z_{j},N)}$ (10)
Since both numerator and denominator are in $\mathbb{Q}$, the result remains strictly rational, effectively computing the probability distribution on a progressively refined rational lattice.
Rational LayerNorm via Newton-Raphson. Layer Normalization requires computing the inverse square root of the variance $\sigma^{2}$, which is irrational. To avoid exiting $\mathbb{Q}$, the EIU employs the Fast Inverse Square Root algorithm adapted for rational numbers using Newton-Raphson iteration. To find $y\approx 1/\sqrt{\sigma^{2}}$, we iterate:
$y_{t+1}=\frac{1}{2}y_{t}\left(3-\sigma^{2}y_{t}^{2}\right)$ (11)
Since operations in this update rule (subtraction, multiplication, scaling by 1/2) are closed under $\mathbb{Q}$, initializing $y_{0}$ as a rational number guarantees that $y_{t}\in\mathbb{Q}$ for all $t$. This method converges quadratically, allowing the EIU to reach a "Reasonable Exactness" threshold (e.g., error $10^{-100}$) within logarithmic steps, far exceeding the precision of FP32 or BF16 without introducing non-deterministic noise.

### v4

@@ Abstract
The prevailing scaling paradigm of Large Language Models (LLMs) rests on a substrate of "Fuzzy" floating-point arithmetic. To mitigate the inherent instability of this approximate foundation, modern architectures have erected a complex scaffolding of structural and numerical heuristics—Complex Residuals, Pre-RMSNorm, Attention Scaling, and Gradient Clipping—consuming significant compute solely to prevent numerical collapse.
We propose a paradigm shift to the "Exact". We introduce the Halo Architecture, grounded in the Rational Field ($\mathbb{Q}$) and powered by a custom Exact Inference Unit (EIU). To resolve the exponential bit-width growth of rational arithmetic, Halo employs a Dual-Ring Topology that unifies two complementary control mechanisms: (1) The Micro-Ring (Continuum Maintenance), which strictly bounds memory complexity via Diophantine Approximation; and (2) The Macro-Ring (Symbolic Alignment), which enforces logical consistency via periodic state collapse.
This stable dual-ring substrate allows for the "Great Dismantling" of numerical scaffolding, reducing the Transformer block to its "Clean" algebraic form (Tabula Rasa). Furthermore, we verify the "Efficiency Paradox": the elimination of gradient noise ($\sigma\to 0$) allows for Macro-Learning Rates, potentially reducing the Total Time-to-Convergence by orders of magnitude. Halo demonstrates that General Intelligence requires the hybridization of continuous fields and discrete chains under a rigorous mathematical framework.
@@ 1 Introduction
We argue that this is due to a fundamental mismatch between the physical substrate (approximate floating-point hardware) and the mathematical requirements of rigorous logic. Logic is discrete and precise; floating-point arithmetic is continuous and approximate. We observe that in deep recursive inference, numerical noise $\epsilon$ accumulates exponentially ($\mathcal{O}(e^{L})$), eventually drowning out the causal signal. We term this phenomenon Semantic Drift. This aligns with recent findings by Thinking Machines Lab [2025], who pinpointed floating-point non-associativity as the culprit for inference instability.
To survive this instability, modern architectures have erected a complex scaffolding of structural and numerical heuristics—Complex Residuals, LayerNorm, Attention Scaling, and Gradient Clipping—consuming significant compute solely to prevent numerical collapse. We term this "Numerical Debt."
@@
However, Halo goes beyond mere exactness. To resolve the tractability of rational arithmetic, we introduce a Dual-Ring Topology. This unifies two complementary control mechanisms: the Micro-Ring (Continuum Maintenance), which utilizes Diophantine Approximation to bound numerical variance with quadratic error decay ($\mathcal{O}(D^{-2})$); and the Macro-Ring (Symbolic Alignment), which periodically collapses the state to enforce logical consistency. This stable substrate allows us to perform the "Great Dismantling"—abolishing all numerical scaffolding to restore the Transformer to its Clean algebraic form.
Contributions.
  1. 1.
    The Exactness Hypothesis: We formally link "Semantic Drift" in deep reasoning to the numerical instability of floating-point arithmetic, proving that $\mathbb{Q}$ is the necessary substrate for infinite-depth logic.
  2. 2.
    The Dual-Ring Topology: We introduce a hierarchical control architecture. The Micro-Ring handles continuous numerical tractability (The Wave), while the Macro-Ring enforces discrete logical consistency (The Particle).
  3. 3.
    The Clean Transformer: Levering the stability of the Dual-Ring, we demonstrate the abolition of "Numerical Band-aids" (LayerNorm, Scaling, Clipping), restoring the Transformer to its pristine algebraic form.
  4. 4.
    The Exact Inference Unit (EIU): We propose a hardware specification featuring CRT-based RNS acceleration and Dual-Modular Redundancy, resolving the efficiency and reliability bottlenecks of exact computation.
@@ 3.2 The Derivative of Truth
We identify two specific mechanical failures in floating-point AD that Halo resolves:
1. The Numerical Event Horizon (Underflow).
In FP16, the smallest positive normal number is $\epsilon_{min}\approx 6\times 10^{-5}$. Any causal dependency spanning long contexts often yields a gradient signal weaker than this threshold, flushing it to zero.
$\nabla_{FP16}\approx 0\quad\text{if }\left|\frac{\partial\mathcal{L}}{\partial w}\right|<\epsilon_{min}$ (3)
This acts as a "Numerical Lobotomy," physically preventing the model from learning subtle, long-range logic.
2. Infinite Sensitivity via Rationals.
In the Halo EIU ($\mathbb{Q}$), there is no Underflow. A gradient of $10^{-100}$ is stored exactly as $(1,10^{100})$. This preserves the direction of descent even when the magnitude is microscopic.
$\nabla_{\mathbb{Q}}\equiv\text{Exact}\quad\forall\text{ magnitude}$ (4)
This allows Halo to optimize through "flat" loss landscapes where FP16 models would stall, justifying the removal of heuristics like Gradient Clipping—since large values now represent true sensitivity, not instability.
@@ 4.1 Algorithmic Formulation
To address the precision-efficiency trade-off, Halo partitions the computational flow into a hierarchical topology consisting of The Light (Stream) and The Rings (Control). We formally define these core components:
  • The Light (Exact Stream): The primary datapath where calculations occur in the arbitrary-precision rational field $\mathbb{Q}$. In this stream, operations are chemically pure: no truncation or rounding occurs ($\epsilon=0$).
  • The Dual Rings: A composite mechanism acting as the "Gatekeeper of Complexity." It consists of the Micro-Ring (Physical Constraint) and the Macro-Ring (Logical Constraint).
4.2 The Macro-Ring: The Rational Semantic Lattice
The Macro-Ring Codebook $\mathcal{C}$ is not merely a set of learned weights but the structural definition of the "Language of Thought" within Halo. To ensure strict adherence to the Exactness Hypothesis and hardware constraints, we design $\mathcal{C}$ as a Rational Semantic Lattice.
1. Algebraic Structure: The Clean Lattice Constraint
Unlike standard VQ-VAE codebooks in $\mathbb{R}^{d}$, the Macro-Ring codebook $\mathcal{C}=\{c_{1},c_{2},\dots,c_{M}\}$ consists of vectors strictly in the rational field $\mathbb{Q}^{d}$. We impose the Clean Lattice Constraint: for every codebook entry $c_{i}$, the complexity of its components is bounded:
$c_{i}\in\mathbb{Q}^{d}\quad\text{s.t.}\quad\forall j,\,\text{Denominator}(c_{i,j})\leq D_{lattice}$ (10)
where $D_{lattice}$ is a fixed small integer (e.g., $2^{16}$). This constraint serves two purposes:
  • Complexity Reset: By forcing the state to collapse onto low-complexity rationals (simple fractions), we guarantee that the subsequent "Light" phase begins with a clean, low-entropy state.
  • Hardware Compatibility: Low denominators ensure that nearest-neighbor search can be performed efficiently using integer arithmetic without overflow.
2. Initialization: Seeding the Semantic Manifold
The topology of the reasoning space is determined by its initialization. We propose:
  • Axiomatic Seeding (Logic/Math): Initialize with "Logical Atoms"—fundamental rational vectors (e.g., $[1,0],[1/2,1/2]$). This provides a strong inductive bias for rigorous deduction.
  • Density-Based Seeding (General): We project pre-trained embeddings onto regions of high probability density in $\mathbb{Q}^{d}$, bridging the gap between "Fuzzy" and "Exact" domains.
3. Training: Rational Vector Quantization (RVQ)
Training involves aligning the continuous rational stream to the discrete lattice $\mathcal{C}$. We utilize a modified update rule that projects gradients back onto the rational field. For a selected codebook vector $c_{i^{*}}$, the update step is:
$c_{i^{*}}\leftarrow\text{RationalApprox}(c_{i^{*}}-\eta\cdot\nabla\mathcal{L},\,D_{lattice})$ (11)
This ensures that the codebook evolves to match the data distribution while remaining strictly within the domain of valid hardware-representable rationals.
4. Operational Use: The Wavefunction Collapse
During inference, the Macro-Ring applies a deterministic "collapse" operator:
@@ 4. Operational Use: The Wavefunction Collapse
This operation is discontinuous and non-linear. It acts as a measurement of the "semantic quantum state," forcing the superposition of possibilities in $H_{\mathbb{Q}}$ to collapse into a definite logical anchor $c_{i^{*}}$.
Table 1: The Core Distinction: Micro-Ring vs. Macro-Ring
Feature Micro-Ring (Continuum Maintenance) Macro-Ring (Logical Anchoring)
Math Object Continuous Points (Farey Sequence $\mathcal{F}_{N}$) Discrete Vectors (Rational Lattice $\mathcal{C}$)
Operation Approximation: Find closest rational $p/q$ Quantization: Find closest semantic vector $c_{i}$
Nature Local, Smooth, Unique Global, Discontinuous, Many-to-One
Role of Truth Numerical Truth: Preserves signal magnitude Logical Truth: Collapses to valid concept
Analogy Noise Filter: Removes high-freq hiss Trap: Locks signal into valid state
Algorithm 1 The Macro-Ring Mechanism: Rational Semantic Lattice
0: Input State $H_{\mathbb{Q}}$, Lattice Codebook $\mathcal{C}=\{c_{1},\dots,c_{M}\}$, Max Denominator $D_{lattice}$
0: Anchored State $H_{anchored}$ (Inference) or Updated Codebook $\mathcal{C}^{\prime}$ (Training)
1:Function LatticeCollapse($H_{\mathbb{Q}}$) {Inference Mode}
2:  // 1. Metric: Find nearest semantic anchor in $\mathbb{Q}^{d}$
3:  $i^{*}\leftarrow\operatorname{argmin}_{i}\|H_{\mathbb{Q}}-c_{i}\|_{\mathbb{Q}}^{2}$
4:  // 2. Collapse: Snap continuous state to discrete lattice
5:  return $c_{i^{*}}$
6:End Function
7:Function LatticeUpdate($H_{\mathbb{Q}},\mathcal{C},\eta$) {Training Mode}
8:  // 1. Standard Gradient Descent Step (in Float)
9:  $i^{*}\leftarrow\operatorname{argmin}_{i}\|H_{\mathbb{Q}}-c_{i}\|^{2}$
10:  $g\leftarrow\nabla_{c_{i^{*}}}\mathcal{L}_{commitment}$
11:  $c_{temp}\leftarrow c_{i^{*}}-\eta\cdot g$
12:  // 2. The Lattice Constraint: Project back to hardware-friendly rationals
13:  $n_{new},d_{new}\leftarrow\operatorname{RationalApprox}(c_{temp},\epsilon_{tol},D_{lattice})$
14:  $c_{i^{*}}^{\prime}\leftarrow\operatorname{Simplify}(n_{new},d_{new})$
15:  return $\mathcal{C}^{\prime}$ with updated $c_{i^{*}}^{\prime}$
16:End Function
4.3 The Micro-Ring: Continuous Tractability
While the Macro-Ring handles semantics, the physical tractability of the infinite-precision stream is governed by the Micro-Ring (evolved from the Halo-N paradigm).
Mechanism: Unlike rigid fixed-point quantization, the Micro-Ring employs a "Natural Saturation" policy. It is triggered only when the denominator’s bit-width exceeds the hardware capacity $B_{max}$ (e.g., 4096 bits). Instead of a learned codebook, it utilizes Diophantine Approximation to project the state onto the Farey Sequence $\mathcal{F}_{N}$. This maintains the Wave Nature—continuous and differentiable—while preventing "Bit-width Explosion."
Algorithm 2 Micro-Ring Mechanism (Diophantine Projector)
1:Function DiophantineProject($H_{\mathbb{Q}}$)
@@ 4.3 The Micro-Ring: Continuous Tractability
3:for all element $h_{ij}$ in tensor $H_{\mathbb{Q}}$ do
4:  $x_{val}\leftarrow h_{ij}.num/h_{ij}.denom$
5:  {Best Rational Approximation via Continued Fractions}
6:  $p,q\leftarrow\text{TruncatedEuclidean}(h_{ij}.num,h_{ij}.denom,D_{max})$
7:  $H_{new}[i,j]\leftarrow\text{Rational}(p,q)$
8:end for
9:return $H_{new}$
10:End Function
4.4 The Dual-Ring Synthesis: The Clean Transformer
In standard Transformers, components like LayerNorm, Attention Scaling ($1/\sqrt{d}$), and Gradient Clipping are strictly necessary to prevent floating-point variance explosion. However, with the Micro-Ring bounding numerical variance and the Macro-Ring enforcing logical rigidity, we perform the "Great Dismantling." We term this final state The Clean Transformer (Tabula Rasa).
4.4.1 The Abolition of Heuristics
Supported by the EIU and the Dual-Ring Protocol, we remove the "Numerical Debt" and restore the network to its pristine form:
  • No LayerNorm (Variance Conservation): Standard LayerNorm forces statistical normalization to prevent variance explosion. In Halo, the Micro-Ring actively manages denominator complexity, ensuring the "Rational Wave" remains bounded. We remove LayerNorm entirely, enabling "Raw Signal" propagation where causal magnitude is preserved.
  • No Attention Scaling (Infinite Dynamic Range): The standard $1/\sqrt{d}$ factor artificially suppresses dot products. The EIU’s 4096-bit dynamic range ($10^{1200}$) accommodates unscaled dot products ($Q\cdot K^{T}$), enabling "Super-Sharp" attention distributions that represent absolute causal certainty.
  • No Gradient Clipping (True Sensitivity): In floating-point, large gradients imply instability (Exploding Gradients). In exact arithmetic, they imply high causal sensitivity. We preserve these strong signals to accelerate convergence.
4.4.2 Algorithm: The Clean Transformer Protocol
Consequently, the Halo Inference Loop is reduced to its algebraic essence. We utilize color-coded annotations to highlight the distinct roles of the dual rings in this "clean" topology.
Algorithm 3 The Clean Transformer Protocol (Dual-Ring)
0: Input $X$, Depth $T$, Macro Interval $K$, Micro Limit $B_{max}$
0: Probability $P(y|X)$
@@ 4.4.2 Algorithm: The Clean Transformer Protocol
2:for $l=1$ to $T$ do
3:  // Phase A: The Light (Original Algebraic Form)
4:  // No LayerNorm. No Scaling. Pure Causality.
5:  $Q,K,V\leftarrow H_{l-1}W_{q},H_{l-1}W_{k},H_{l-1}W_{v}$
6:  $A\leftarrow\text{RatSoftmax}(Q\cdot K^{T})$ {Unscaled Attention}
7:  $H_{attn}\leftarrow A\cdot V$
8:  $H_{mlp}\leftarrow\text{RationalFFN}(H_{attn})$
9:  $H_{temp}\leftarrow H_{mlp}+H_{l-1}$ {Clean Residual Path}
10:  // Phase B: Micro-Ring (Continuum Maintenance)
11:  // Replaces LayerNorm via Diophantine Projection
12:  if $\text{CheckSaturation}(H_{temp},B_{max})$ then
13:   $H_{temp}\leftarrow\text{DiophantineProject}(H_{temp})$
14:  end if
15:  // Phase C: Macro-Ring (Logical Anchoring)
16:  // Replaces Clipping via Symbolic Alignment
17:  if $l\pmod{K}==0$ then
18:   $H_{l}\leftarrow\text{SymbolicAlign}(H_{temp})$
19:  else
20:   $H_{l}\leftarrow H_{temp}$
21:  end if
22:end for
23:return $\text{Softmax}(\text{ToFloat}(H_{T})\cdot W_{vocab})$
4.4.3 Modes of Cognition: Continuous vs. Chain Reasoning
The Dual-Ring topology enables Halo to unify two previously distinct modes of cognition, mirroring the Wave-Particle Duality:
  1. 1.
    Continuous Reasoning (The Wave): Between Macro-Ring resets (steps $l\neq nK$), the model operates in "The Light." Here, reasoning is a continuous fluid. The model maintains a superposition of possibilities without being forced to choose a discrete symbol. This corresponds to Intuition or "System 1"—fast, high-dimensional, and nuanced.
  2. 2.
    Chain Reasoning (The Particle): At Macro-Ring boundaries (steps $l=nK$), the model undergoes Symbolic Alignment. The wave collapses into a discrete state. This forces the model to "make a decision," anchoring its thought process to a verifiable logical checkpoint. This corresponds to Chain-of-Thought or "System 2"—discrete, sequential, and rigorous.
By oscillating between these states—Continuous Reasoning to gather information and Chain Reasoning to define truth—Halo achieves an Infinite-Depth Reasoning capability that is both creative (Wave) and correct (Particle).
@@ 4.5 Implementation: The Exact Inference Unit (EIU)
    The CRT Duality (Parallelism & Integrity): We leverage the algebraic properties of the Chinese Remainder Theorem (CRT) to achieve both speed and safety.
    • For Acceleration (RNS): The EIU decomposes high-precision integers into independent residues, transforming monolithic $\mathcal{O}(N^{2})$ multiplication into $\mathcal{O}(1)$ parallel MIMD operations (see Appendix APPENDIX F: OPERATIONAL ACCELERATION VIA RESIDUE NUMBER SYSTEMS).
    • For Reliability (SDC): It simultaneously exploits this modular structure to enforce Dual-Modular Redundancy via Mersenne primes, preventing Silent Data Corruption without stalling the pipeline (see Appendix C).
4.6 Constructive Rationality: Handling Non-Linearities
A theoretical challenge for Rational Arithmetic is that standard activation functions (e.g., Softmax, GeLU) rely on transcendental operations ($e^{x}$) that yield irrational numbers, theoretically exiting the field $\mathbb{Q}$. In the Halo architecture, we resolve this through Elimination and Construction:
1. Elimination of Irrational Norms.
Standard LayerNorm requires computing $\sigma=\sqrt{Var(x)}$, an irrational operation demanding iterative approximation (e.g., Newton-Raphson). By abolishing LayerNorm (Section 4.4), Halo eliminates the primary source of irrationality in the Transformer block, significantly reducing the EIU’s complexity.
2. Constructive Softmax (Rational Taylor Series).
For the attention mechanism ($A=\text{softmax}(QK^{T})$), we replace the transcendental exponential $e^{x}$ with its constructive definition via Taylor Series:
$\text{RatExp}(x,N)=\sum_{k=0}^{N}\frac{x^{k}}{k!}$ (13)
Since $x\in\mathbb{Q}$, every term in the series remains strictly rational. Unlike floating-point libraries that rely on look-up tables with fixed error, the EIU dynamically adjusts $N$ to match the precision of the datapath. This ensures that Softmax remains a closed operation within $\mathbb{Q}$, preserving the "Exactness Hypothesis" end-to-end.